During exercise glucose consumption increases to 60 :mol min-1 100 g-1. Assume arterial [glucose]
does not change. Blood flow increases to 0.60 mL min-1 g-1.
G. What is the venous concentration of glucose during exercise?
Answer #1

According to the given data

Here QM = QV ( [Glucose]artery - [Glucose]vein)
0.60 x 10-6 mol min-1 g-1 = 0.60 mL min-1 g-1 ( 4.99 x 10-6 mol mL-1 - X)
Solving for [Glucose]vein = 4.99 x 10-6 mol mL-1 - 1 x 10-6 mol mL-1 = 3.99 x 10-6 mol mL-1
= 3.99 mM
= 3.99 x 10-3 mol L-1 x 180.2 g mol-1 x 0.1 L dL-1 = 72.9 mg%